True or False

6. Carbon-14 is a radioactive isotope of carbon.   .

7. The 3 isotopes of hydrogen differ in the number of protons?

8. The kinetic energy does a 1000 Kg car have going 15 m/s is 15000 Joules.

9. A metal weighs 100 kg and has a volume of 4 m3 has a density400 kg/m.

10. Radiation is the process of transferring heat from one material to another by contact.

11 The normal hearing range is between 20 hz and 80,000 hz
Answer #1

6.

True, C-14 is radioactive

7.

False, the 3 isotopes have same number of proton, which is 1 proton in each of them.

8.

KE = 0.5*m*V^2 = 0.5*1000*15^2 = 112500 J

False

9.

density = mass/Volume = 100/4 = 25 kg/m^3

False

10.

True.

11.

Normal hearing range is 20 Hz to 20000 Hz

False

Please Upvote